OpenStudy (anonymous):

Determine which polynomial is a difference of two squares. a2 + 36 a2 − 36 a2 − 12 a2 + 12

jimthompson5910 (jim_thompson5910):

Hint: a "difference" is the result of a subtraction problem Ex: 12 - 5 is a difference while 12+5 is a sum.

OpenStudy (ghazi):

b, just see 36 is square of 6 and a^2 is square of a

OpenStudy (dariyana):

b

jimthompson5910 (jim_thompson5910):

A square is a number y such that y = x^2 for some integer x

OpenStudy (ghazi):

and take their difference

OpenStudy (dariyana):

a^2-36=(a-6)(a+6)